# Seg2Any: Open-set Segmentation-Mask-to-Image Generation with Precise Shape and Semantic Control

### v1

@@ C.1 Comparison with Existing Dense Caption Datasets
SACap-1M (ours) SA-1B [15] 1M Qwen2-VL-72B [35] 14.1 5.7
C.2 Data annotation pipeline
@@
\@slowromancap
i@) Image Filtering. Images are filtered based on size and aspect ratio, keeping those with width and height between 1000 and 3000 pixels and aspect ratios from 0.6 to 1.8. To further ensure visual quality, images with a LAION-Aesthetics [32] score below $5$ are excluded.
\@slowromancap
ii@) Entity Extraction. While SA-1B provides numerous accurate, class-agnostic segmentation masks—averaging over 100 per image—many are nested and redundant. To mitigate this redundancy, we retain only the top-level masks, which generally correspond to the primary objects in the image, and discard any masks fully enclosed by others (as shown in Figure 8). Additionally, masks occupying less than $1\%$ of the image area are considered too small and removed. Finally, only images containing between 1 and 20 valid masks are retained for further annotation.
\@slowromancap
iii@) Regional and Global Caption Annotation. We utilize the vision-language model Qwen2-VL-72B [35] for both regional and global captioning. For regional captioning, we incorporate the bounding box coordinates of each segment into the text prompt (see Figure 8), enabling the model to produce accurate and context-aware descriptions of each local region. This approach yields detailed annotations, with an average of 14.1 words per entity and 58.6 words per image, supporting fine-grained segmentation-to-image generation.
@@
C.3 SACap-Eval Benchmark
@@
Figure 10: Examples from the SACap-1M dataset.
Appendix D More Qualitative Results
Figure 11: More qualitative results using the Seg2Any method.
We provide additional qualitative results in Figure 11 to illustrate the effectiveness of our approach. Seg2Any shows strong capability in generating high-quality images that faithfully adhere to detailed entity descriptions, enabling flexible and precise control over complex visual attributes such as color, texture, and shape.
Appendix E Broader Impacts
The proposed method for segmentation-mask-to-image synthesis has potential applications in creative industries, design, and education. However, like other generative models, its misuse could result in the production of misleading or inappropriate content. Our approach may also inherit biases present in the training datasets, potentially reinforcing certain stereotypes. Responsible usage and further investigation into fairness and robustness are important, but a comprehensive analysis is beyond the scope of this work.

### v3

@@ E.1 Comparison with Existing Dense Caption Datasets
Pix2Cap-COCO [50] COCO [23] 20k GPT-4V [1] and human correction 22.94 8.14
SACap-1M (ours) SA-1B [16] 1M Qwen2-VL-72B [43] 14.1 5.9
E.2 Data annotation pipeline
@@
I) Image Filtering. Images are filtered based on size and aspect ratio, keeping those with width and height between 1000 and 3000 pixels and aspect ratios from 0.6 to 1.8. To further ensure visual quality, images with a LAION-Aesthetics [35] score below $5$ are excluded.
II) Entity Extraction. While SA-1B provides numerous accurate, class-agnostic segmentation masks—averaging over 100 per image—many are nested and redundant. To mitigate this redundancy, we retain only the top-level masks, which generally correspond to the primary objects in the image, and discard any masks fully enclosed by others (as shown in Figure 12). Additionally, masks occupying less than $1\%$ of the image area are considered too small and removed. Finally, only images containing between 1 and 20 valid masks are retained for further annotation.
III) Regional and Global Caption Annotation. We utilize the vision-language model Qwen2-VL-72B [43] for both regional and global captioning. For regional captioning, we incorporate the bounding box coordinates of each segment into the text prompt (see Figure 12), enabling the model to produce accurate and context-aware descriptions of each local region. This approach yields detailed annotations, with an average of 14.1 words per entity and 58.6 words per image, supporting fine-grained segmentation-to-image generation.
@@
Figure 14: Examples from the SACap-1M dataset.
E.3 SACap-Eval Benchmark
@@
Appendix F Broader Impacts
The proposed method for segmentation-mask-to-image synthesis has potential applications in controllable image generation [21, 14], video generation [40, 41, 39], and the construction of image segmentation datasets. However, like other generative models, its misuse could result in the production of misleading or inappropriate content. Our approach may also inherit biases present in the training datasets, potentially reinforcing certain stereotypes. Responsible usage and further investigation into fairness and robustness are important, but a comprehensive analysis is beyond the scope of this work.